# The maximal correlation coefficient associated with the minimum

### v1

@@ 1 Introduction
By [BS25, Corollary 2.2], the upper bounds in (1) and (2) are sharp when each $X_{i}$ follows an exponential distribution. We thus wonder whether they are also sharp for other distributions, see the second open problem in [CC24, Section 6].
@@
$R=\frac{\prod_{i=1}^{n}p_{i}-\prod_{i=1}^{m}p_{i}\prod_{i=\ell+1}^{n}p_{i}}{\sqrt{\prod_{i=1}^{m}p_{i}(1-\prod_{i=1}^{m}p_{i})\prod_{i=\ell+1}^{n}p_{i}(1-\prod_{i=\ell+1}^{n}p_{i})}}.$ (3)
@@
$=\frac{\prod_{i=1}^{n}(1-p_{i})-\prod_{i=1}^{m}(1-p_{i})\prod_{i=\ell+1}^{n}(1-p_{i})}{\sqrt{\prod_{i=1}^{m}(1-p_{i})(1-\prod_{i=1}^{m}(1-p_{i}))\prod_{i=\ell+1}^{n}(1-p_{i})(1-\prod_{i=\ell+1}^{n}(1-p_{i}))}}.$
@@ 2 Preliminary
Suppose $(X_{k},Y_{k})$ converges to $(X,Y)$ in law. By [CC24, Lemma 2.3],
@@
Consider a finite or countably infinite collection $(X_{i},Y_{i})_{i\in I}$ of independent random tuples. Then we have that
$R((X_{i})_{i\in I},(Y_{i})_{i\in I})=\sup_{i\in I}R(X_{i},Y_{i}).$
If the collection is finite, it is known as the Csáki-Fischer identity; see [CF63, Theorem 6.2] and [Wit75, Theorem 1]. For the case $|I|=\infty$, similar to the proof of [CC24, Theorem 1.1], we can prove Lemma 2.1 as a consequence of the Csáki-Fischer identity and [CC24, Lemma 2.3].
@@ 3 Continuous distribution
By [BS25, Corollary 2.2], the upper bound in (2) is sharp when each $X_{i}$ follows an exponential distribution. For general continuous random variables $X_{1},X_{2},\ldots,X_{n}$ with the common distribution function $F(x)$, define
$Y_{i}=\varphi(X_{i}),$
where $\varphi(x)=-\ln(1-F(x))$. Note that $\varphi$ is a non-decreasing function. Moreover, $Y_{1},Y_{2},\ldots,Y_{n}$ are i.i.d. exponential random variables with parameter 11. Hence, we have that
@@
The proof is complete by combining the lower bound (3) with the upper bound (2).
@@
Alternatively, the lower bound (3) could be derived from Theorem 1.3. Indeed, let
$h(x)=1_{(\delta(p),\infty)}(x),$
where $P(X_{i}>\delta(p))=p$. Then
@@ 4.1 Bernoulli distribution
Consider a random vector $\left(\xi,\eta\right)$ with the joint distribution in Table 1.
$\xi\backslash\eta$ $c$ $d$
$a$ $p_{ac}$ $p_{ad}$
$b$ $p_{bc}$ $p_{bd}$
@@
$R(\xi,\eta)=\left.\left|\det\begin{pmatrix}p_{ac}&p_{ad}\\ p_{bc}&p_{bd}\end{pmatrix}\right.\right|/\sqrt{p_{a}p_{b}p_{c}p_{d}},$
where $p_{a}=P(\xi=a)$, $p_{b}=P(\xi=b)$, $p_{c}=P(\eta=c)$ and $p_{d}=P(\eta=d)$.
@@
By Lemma 4.1, we have that
$R(X,Y)=\frac{\prod_{i=1}^{n}p_{i}-\prod_{i=1}^{m}p_{i}\prod_{i=\ell+1}^{n}p_{i}}{\sqrt{\prod_{i=1}^{m}p_{i}(1-\prod_{i=1}^{m}p_{i})\prod_{i=\ell+1}^{n}p_{i}(1-\prod_{i=\ell+1}^{n}p_{i})}}.$
Remark 4.1.
Recall the definition of $R_{m,\ell}(p)$ in (4). Note that $R_{m,\ell}(p)$ is continuous on $(0,1)$ with $R_{m,\ell}(0+)=0$ and $R_{m,\ell}(1-)=(m-\ell)/\sqrt{m(n-\ell)}$. Moreover, $R_{m,\ell}(p)$ is non-decreasing.
Proof.
From (4), it is not difficult to check the continuity of $p\mapsto R_{m,\ell}(p)$ and to obtain the values of $R_{m,\ell}(0+)$ and $R_{m,\ell}(1-)$. For the monotonicity of $R_{m,\ell}$, let $a=(n+\ell-m)/2$, $b=m$-$\ell$ and $c=n$-$\ell$. Then for $0<p<1$,
@@
For $x>0$, define $f(x)=x/(1-p^{x})$. Then
$f^{\prime}(x)=\frac{1-p^{x}+xp^{x}\ln p}{(1-p^{x})^{2}}>0.$
Hence, $f(x)$ is increasing. Since $b\leq m$, we have that $f(m)-f(b)\geq 0$. Similarly, $f(c)-f(b)\geq 0$. Therefore, we have that
@@ 4.2 Geometric distribution
Let $U=\min_{i:1\leq i\leq m}X_{i}$ and $V=\min_{j:\ell+1\leq j\leq n}X_{j}$. Then
$U=\min\{k\geq 1:U_{k}=0\},$
where
$U_{k}=\min_{i:1\leq i\leq m}B_{ik}.$
Similarly,
$V=\min\{k\geq 1:V_{k}=0\},$
where
$V_{k}=\min_{j:\ell+1\leq j\leq n}B_{jk}.$
On the other hand,
$U_{1}=1_{U\geq 2}\text{ and }V_{1}=1_{V\geq 2}.$
@@
$R(U_{k},V_{k})=R(U_{1},V_{1})$
$=\frac{\prod_{i=1}^{n}(1-p_{i})-\prod_{i=1}^{m}(1-p_{i})\prod_{i=\ell+1}^{n}(1-p_{i})}{\sqrt{\prod_{i=1}^{m}(1-p_{i})(1-\prod_{i=1}^{m}(1-p_{i}))\prod_{i=\ell+1}^{n}(1-p_{i})(1-\prod_{i=\ell+1}^{n}(1-p_{i}))}}.$
@@
Finally, since $U$ is a measurable function of $(U_{k})_{k\geq 1}$ and $V$ is a measurable function of $(V_{k})_{k\geq 1}$, we have that
$R(U,V)\leq R((U_{k})_{k\geq 1},(V_{k})_{k\geq 1}).$
On the other hand, since $U_{1}$ is a measurable function of $U$ and $V_{1}$ is a measurable function of $V$, we have that
$R(U,V)\geq R(U_{1},V_{1}).$
Finally, we conclude that
$R(U,V)=R(U_{1},V_{1})$
$=\frac{\prod_{i=1}^{n}(1-p_{i})-\prod_{i=1}^{m}(1-p_{i})\prod_{i=\ell+1}^{n}(1-p_{i})}{\sqrt{\prod_{i=1}^{m}(1-p_{i})(1-\prod_{i=1}^{m}(1-p_{i}))\prod_{i=\ell+1}^{n}(1-p_{i})(1-\prod_{i=\ell+1}^{n}(1-p_{i}))}}.$
@@ 4.3 Binomial distribution
We need the following lemma on the monotonicity of $k\mapsto p(k)$.
@@
$\Delta:=\left(P(X_{i}\geq k)\right)^{2}-P(X_{i}\geq k+1)P(X_{i}\geq k-1)\geq 0$
for $k\geq 1$. Note that
$P(X_{i}\geq k)=P(X_{i}=k)+P(X_{i}\geq k+1),$
$P(X_{i}\geq k-1)=P(X_{i}=k-1)+P(X_{i}\geq k).$
Then we have that
@@
Let us continue with the proof of Theorem 1.5. We have the following equality in distribution:
$X_{i}\overset{(d)}{=}\min\{k\geq 1:B_{ik}=0\}-1.$
Let $U=\min_{i:1\leq i\leq m}X_{i}$ and $V=\min_{j:\ell+1\leq j\leq n}X_{j}$. Then
$U=\min\{k\geq 1:U_{k}=0\}-1,$
where
$U_{k}=\min_{i:1\leq i\leq m}B_{ik}.$
Similarly,
$V=\min\{k\geq 1:V_{k}=0\}-1,$
where
$V_{k}=\min_{j:\ell+1\leq j\leq n}B_{jk}.$
On the other hand,
$U_{1}=1_{U>0}\text{ and }V_{1}=1_{V>0}.$
Since $U_{1}$ is a measurable function of $U$ and $V_{1}$ is a measurable function of $V$, we have that
$R(U_{1},V_{1})\leq R(U,V).$
Since $U$ is a measurable function of $(U_{k})_{k\geq 1}$ and $V$ is a measurable function of $(V_{k})_{k\geq 1}$, we have that
$R(U,V)\leq R((U_{k})_{k\geq 1},(V_{k})_{k\geq 1}).$
By Lemma 2.1, Theorem 1.3, Remark 4.1 and Lemma 4.2,

### v2

@@ 1 Introduction
When each $X_{i}$ follows an exponential distribution with parameter $\lambda$, we find that
$=R\left(\min\left\{\min_{i:1\leq i\leq\ell}X_{i},\min_{i:\ell+1\leq i\leq m}X_{i}\right\},\min\left\{\min_{j:\ell+1\leq j\leq m}X_{j},\min_{j:m+1\leq j\leq n}X_{j}\right\}\right)$
$=R\left(\min(V_{1},V_{2}),\min(V_{2},V_{3})\right),$
where $V_{1},V_{2},V_{3}$ are independent exponential random variables with parameters $\lambda\ell$, $\lambda(m-\ell)$, and $\lambda(n-m)$, respectively. Then by [BS25, Corollary 2.2], the upper bounds in (1) and (2) are sharp when each $X_{i}$ follows an exponential distribution. We thus wonder whether they are also sharp for other distributions, see the second open problem in [CC24, Section 6].
@@
$R=\frac{\prod_{i=1}^{n}p_{i}-\left(\prod_{i=1}^{m}p_{i}\right)\left(\prod_{i=\ell+1}^{n}p_{i}\right)}{\sqrt{\left(\prod_{i=1}^{m}p_{i}\right)\left(1-\prod_{i=1}^{m}p_{i}\right)\left(\prod_{i=\ell+1}^{n}p_{i}\right)\left(1-\prod_{i=\ell+1}^{n}p_{i}\right)}}.$ (3)
@@
$R\left(\min_{i:1\leq i\leq m}X_{i},\min_{j:\ell+1\leq j\leq n}X_{j}\right)=$
$\frac{\prod_{i=1}^{n}\left(1-p_{i}\right)-\left(\prod_{i=1}^{m}\left(1-p_{i}\right)\right)\left(\prod_{i=\ell+1}^{n}\left(1-p_{i}\right)\right)}{\sqrt{\left(\prod_{i=1}^{m}\left(1-p_{i}\right)\right)\left(1-\prod_{i=1}^{m}\left(1-p_{i}\right)\right)\left(\prod_{i=\ell+1}^{n}\left(1-p_{i}\right)\right)\left(1-\prod_{i=\ell+1}^{n}\left(1-p_{i}\right)\right)}}.$
@@ 2 Preliminary
Suppose that $(X_{k},Y_{k})$ converges to $(X,Y)$ in law in some proper space. By [CC24, Lemma 2.3],
@@
Consider a sequence of independent random variables $(X_{i},Y_{i})_{i\geq 1}$, and suppose that $X_{i}$ takes values in some Polish space $\mathcal{X}_{i}$, while $Y_{i}$ takes values in some Polish space $\mathcal{Y}_{i}$. Then we have that
$R((X_{i})_{i\geq 1},(Y_{i})_{i\geq 1})=\sup_{i\geq 1}R(X_{i},Y_{i}).$
Proof.
Suppose that $d_{X,i}(\cdot,\cdot)$ is a metric on the Polish space $\mathcal{X}_{i}$, and $d_{Y,i}(\cdot,\cdot)$ is a metric on the Polish space $\mathcal{Y}_{i}$. Then we define a metric $d_{X}(\cdot,\cdot)$ on the product space $\mathcal{X}=\prod_{i\geq 1}\mathcal{X}_{i}$: For $x=(x_{i})_{i\geq 1}$ and $\tilde{x}=(\tilde{x}_{i})_{i\geq 1}$, we define
$d_{X}(x,\tilde{x})=\sum_{i=1}^{\infty}\min(d_{X,i}(x_{i},\tilde{x}_{i}),1)/2^{i}.$
Then $(\mathcal{X},d_{X})$ is a product Polish space. Similarly, we construct the product Polish space $(\mathcal{Y},d_{Y})$. Then for each $i\geq 1$, we fix a particular point $x_{i}^{0}$ in the Polish space $\mathcal{X}_{i}$ and a particular point $y_{i}^{0}$ in the Polish space $\mathcal{Y}_{i}$. Let $((X_{i})_{i\geq 1},(Y_{i})_{i\geq 1})$ have the joint distribution $\mu$ on the product Polish space $\mathcal{X}\times\mathcal{Y}$. For $m\geq 1$, define
$\bm{X}^{(m)}=(X_{1},X_{2},\cdots,X_{m},x_{m+1}^{0},x_{m+2}^{0},\ldots)\text{ and }\bm{Y}^{(m)}=(Y_{1},Y_{2},\cdots,Y_{m},y_{m+1}^{0},y_{m+2}^{0},\ldots),$
with the joint distribution $\mu_{m}$. Then by Csáki-Fischer identity (see [CF63, Theorem 6.2] or [Wit75, Theorem 1]),
$R(\mu_{m})=R\left(\bm{X}^{(m)},\bm{Y}^{(m)}\right)=R\left((X_{i})_{1\leq i\leq m},(Y_{i})_{1\leq i\leq m}\right)=\max_{1\leq i\leq m}R(X_{i},Y_{i}).$
As $m\to\infty$, $\left(\bm{X}^{(m)},\bm{Y}^{(m)}\right)$ converges almost surely to $((X_{i})_{i\geq 1},(Y_{i})_{i\geq 1})$. Hence, as $m\to\infty$, the joint distribution $\mu_{m}$ of $\left(\bm{X}^{(m)},\bm{Y}^{(m)}\right)$ converges weakly to the joint distribution $\mu$ of $((X_{i})_{i\geq 1},(Y_{i})_{i\geq 1})$. By (7) we have that,
$R\left((X_{i})_{i\geq 1},(Y_{i})_{i\geq 1}\right)\leq\liminf_{m\to\infty}R\left(\bm{X}^{(m)},\bm{Y}^{(m)}\right)=\sup_{i\geq 1}R(X_{i},Y_{i}).$
On the other hand, for each $i\geq 1$, $X_{i}$ is a measurable function of $(X_{j})_{j\geq 1}$ and $Y_{i}$ is a measurable function of $(Y_{j})_{j\geq 1}$. Hence, we have that
$R(X_{i},Y_{i})\leq R\left((X_{j})_{j\geq 1},(Y_{j})_{j\geq 1}\right).$
Then we have the conclusion. ∎
@@ 3 Continuous distribution
As explained in the introduction, the upper bound in (2) is sharp when each $X_{i}$ follows an exponential distribution. For general continuous random variables $X_{1},X_{2},\ldots,X_{n}$ with the common distribution function $F(x)$, define $Y_{i}=\varphi(X_{i})$, where $\varphi(x)=-\ln(1-F(x))$. Note that $\varphi$ is a non-decreasing function. Moreover, $Y_{1},Y_{2},\ldots,Y_{n}$ are i.i.d. exponential random variables with parameter $1$. Hence, we have that
@@
By (2) and (3), the proof is complete.
@@
Alternatively, the lower bound (3) could be derived from Theorem 1.3. Indeed, let $h(x)=1_{(\delta(p),\infty)}(x)$, where $P(X_{i}>\delta(p))=p$. Then
@@ 4.1 Bernoulli distribution
Consider a random vector $\left(\xi,\eta\right)$ taking values in $\{0,1\}^{2}$, with the joint distribution in Table 1.
$\xi\backslash\eta$ 0 11
0 $p_{00}$ $p_{01}$
11 $p_{10}$ $p_{11}$
@@
$R(\xi,\eta)=\left.\left|\det\begin{pmatrix}p_{00}&p_{01}\\ p_{10}&p_{11}\end{pmatrix}\right.\right|/\sqrt{(p_{00}+p_{01})(p_{10}+p_{11})(p_{00}+p_{10})(p_{01}+p_{11})}.$
@@
By Lemma 4.1, we have the (3). Finally, we prove the the monotonicity of $p\mapsto R_{m,\ell}(p)$. let $a=(n+\ell-m)/2$, $b=m$-$\ell$ and $c=n$-$\ell$. Then for $0<p<1$,
@@
For $x>0$, define $f(x)=x/(1-p^{x})$. Then $f^{\prime}(x)=\frac{1-p^{x}+xp^{x}\ln p}{(1-p^{x})^{2}}>0$. Hence, $f(x)$ is increasing. Since $b\leq m$, we have that $f(m)-f(b)\geq 0$. Similarly, $f(c)-f(b)\geq 0$. Therefore, we have that
@@ 4.2 Geometric distribution
Let $U=\min_{i:1\leq i\leq m}X_{i}$, $V=\min_{j:\ell+1\leq j\leq n}X_{j}$, and $U_{k}=\min_{i:1\leq i\leq m}B_{ik}$, we have that $U=\min\{k\geq 1:U_{k}=0\}$. Similarly, let $V_{k}=\min_{j:\ell+1\leq j\leq n}B_{jk}$, we have that $V=\min\{k\geq 1:V_{k}=0\}$. On the other hand, $U_{1}=1_{U\geq 2}\text{ and }V_{1}=1_{V\geq 2}$.
@@
$R(U_{k},V_{k})=R(U_{1},V_{1})=$
$\frac{\prod_{i=1}^{n}\left(1-p_{i}\right)-\left(\prod_{i=1}^{m}\left(1-p_{i}\right)\right)\left(\prod_{i=\ell+1}^{n}(1-p_{i})\right)}{\sqrt{\left(\prod_{i=1}^{m}\left(1-p_{i}\right)\right)\left(1-\prod_{i=1}^{m}\left(1-p_{i}\right)\right)\left(\prod_{i=\ell+1}^{n}(1-p_{i})\right)\left(1-\prod_{i=\ell+1}^{n}\left(1-p_{i}\right)\right)}}.$
@@
Finally, since $U$ is a measurable function of $(U_{k})_{k\geq 1}$ and $V$ is a measurable function of $(V_{k})_{k\geq 1}$, we have that $R(U,V)\leq R((U_{k})_{k\geq 1},(V_{k})_{k\geq 1})$. On the other hand, since $U_{1}$ is a measurable function of $U$ and $V_{1}$ is a measurable function of $V$, we have that $R(U,V)\geq R(U_{1},V_{1})$. Finally, we have that $R(U,V)=R(U_{1},V_{1})$ and the proof is complete.
@@ 4.3 Binomial distribution
For $k\geq d+1$ and $i\geq 1$, $B_{ik}=0$. Then we need the following lemma on the monotonicity of $k\mapsto p(k)$.
@@
$\Delta:=\left(P(X_{i}\geq k)\right)^{2}-P(X_{i}\geq k+1)P(X_{i}\geq k-1)\geq 0\text{ for }k\geq 1.$
Note that
@@
Let us continue with the proof of Theorem 1.5. We have the following equality in distribution: $X_{i}\overset{(d)}{=}\min\{k\geq 1:B_{ik}=0\}-1$. Without loss of generality, we assume that the above equalities hold almost surely. Let $U=\min_{i:1\leq i\leq m}X_{i}$, $V=\min_{j:\ell+1\leq j\leq n}X_{j}$, and let $U_{k}=\min_{i:1\leq i\leq m}B_{ik}$, we have that $U=\min\{k\geq 1:U_{k}=0\}-1$. Similarly, let $V_{k}=\min_{j:\ell+1\leq j\leq n}B_{jk}$, we have that $V=\min\{k\geq 1:V_{k}=0\}-1$. On the other hand, $U_{1}=1_{U>0}\text{ and }V_{1}=1_{V>0}$.
Since $U_{1}$ is a measurable function of $U$ and $V_{1}$ is a measurable function of $V$, we have that $R(U_{1},V_{1})\leq R(U,V)$. Since $U$ is a measurable function of $(U_{k})_{k\geq 1}$ and $V$ is a measurable function of $(V_{k})_{k\geq 1}$, we have that $R(U,V)\leq R((U_{k})_{k\geq 1},(V_{k})_{k\geq 1})$. By Lemma 2.1, Theorem 1.3, and Lemma 4.2,